Casas Modernas of Juarez, Mexico, is contemplating a major change in its cost structure. Currently, all of its drafting work- Your answer is partially correct. Calculate the increase in Net income for each alternative if sales increased by $149,000.
Answer #1

Conrtibution Margin ratio = Contribution MArgin / sales Revenue

Manual System = 348000/1740000 = 20%

Computerised System = 1044000/1740000 =60%

Increase in net income = CMRatio * increase in Sales

Manual System = $149,000*20% = $29,800(Answer)

Computerised system = $149,000*60% =$89,400 (Answer)